OpenStudy (anonymous):

what is the following expression equal _____ √ 62-2i^2 _____ A.)√60i B.)8i C.) 8 ____ ____ D.) √62 -√ 2i

OpenStudy (amistre64):

sqrt(64) = 8

OpenStudy (amistre64):

62 - 2i^2 = 62 + 2 = 64

OpenStudy (anonymous):

so would the answer be A

OpenStudy (amistre64):

i^2 = -1; the answer = 8; C

OpenStudy (anonymous):

ohh so C

OpenStudy (anonymous):

okay thanks

OpenStudy (amistre64):

yore welcome :)